Consider arterial blood viscosity µ = .0 025 poise. If the length of the artery is 5.1 cm, radius
3
8 10−
× cm. and 3
1 2 P = P − P = 4×10 dynes/ 2
cm then find (i) maximum peak velocity of
blood and (ii) the shear stress at the wall.
1
Expert's answer
2020-03-25T12:36:10-0400

Here, viscosity of blood is given as "\mu=0.0025 poise" , length of capillary = 5.1 cm, radius="3\times10^{-8}" cm and "P_1-P_2= 4\times10^{12} \frac{dyne}{cm^2}"

we know that maximum velocity in case of fluid flow


Vmax="\frac{\Delta P r^2}{4 \mu l}= \frac{4\times10^{12}\times9\times10^{-16}}{4\times0.0025\times5.1}" = 235 "cm\/s"


and for shear stress we know that,

"\tau=\frac{\Delta P r}{2l}= \frac{4\times10^{12}\times3\times10^{-8}}{2\times5.1}=1.17 \times 10^4 dyne/cm^2"